Journalist Report October 22nd

Mars Crew 197 – Sol 1 – It’s a Wrap (Already??)

Marge Lipton – Crew 197 Journalist

We arrived at dusk last night and were treated to the soft luscious colors of a southwest sunset. Our crew faced earthly challenges like delayed flights and lost luggage, but persevered, fitting 5 people and too many pieces of luggage to count in a Nissan Pathfinder for the two and a half hour drive to the station. We squeezed everything in, but as newly minted Martians this became a bonding experience. Because it’s clear that besides all the engineering and technology in getting humans to Mars, it’s impossible not to take into consideration how best to get a diverse group of people to get along in such close quarters.

Our first day, today was the start of the MDRS season and involved a lot of housecleaning, from mopping and vacuuming, to setting up new posts to build the simulated pressurized tunnel allowing us to walk between different buildings. After the posts were pounded into the ground, we then WRAPPED the tarps overhead. Rovers had to be moved into position and the MDRS car needed to be taken in for a tune-up.

Crew members discussed what each wanted to accomplish on this mission, and tried to figure out especially, how we were going to remain in sim while also trying to get the best VR coverage of the sites James and Shannon wanted. In practice, this meant, to break sim or not wear a helmet or not on the EVA.

We were interrupted in our WRAPPING of the tunnel by rain, later followed by thunderstorms. So we took the opportunity to make WRAPS of tuna, avocado and tomato.

So as I said, that’s a wrap for today.

Operations Report October 22nd

Operations Report 22OCT2018

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Batteries for power system

Notes on non-nominal systems: Batteries arrive tomorrow and will be installed on Friday!

Generator (hours run): 24

Solar— SOC n/a %

Diesel – 80 %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (axillary tank) – 0 gallons.

Water (static) – 500 gallons

GreenHab water tank: 200 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda used to move Deimos into charging position, the 350 that was not starting needed fuel, because I switched to the reserve tank and it worked fine. This is interesting, because it really did not look empty, it looked about 1/3 fuel. All ATVs, if used, should have a full tank when leaving MDRS. The three 350’s and the 300 have been taken out of sim service. The Honda remains, but only to be used if a rover breaks down and needs tows back to MDRS. Crews will not be trained on the use of ATVs, but one experienced person or chosen person per crew will be trained on how to tow.

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs:

Deimos rover used: Yes, towed into position in the charging bay next to Hab. I will work on it there,


Beginning charge:

Ending charge:

Currently charging: Yes

Sojourner rover used: yes


Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: no.


Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: no


Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: no


Beginning charge:

Ending charge:

Currently charging: no

HabCar used and why, where? Was used in the am to haul water, then driven to Bicknell for the following repairs: New windshield, new taillight, oil change and safety inspection. We will pick it up at the end of the week and get new tabs for it.

Crew Transport Vehicle: Yes, to drop the HabCar off and to get fence posts for the tunnel

General notes and comments: Work on the tunnel was started today but not finished due to (you guessed it!) rain. Most of the fenceposts were reset and four of the tarps are in place.

Summary of internet: All three accounts nominal. HAL is
operational. More information will be coming about this system.

Summary of suits and radios: Suits have arrived and are being charged for training in the morning.

Summary of Hab operations: All materials have been removed from in front of the Hab except the rocks, which we are referring to as “the in situ resource material”. I will move it over time.

There is a leak in the drain pipe from the summer work done on the plumbing. It is where the old science sink used to drain. For now, we are using the upstairs sink normally, and catching the water in a bucket using the contractor can fix it.

We will be turning on the hot water in the morning and conducting a propane test in the morning.

Summary of GreenHab operations: We began using the big blue water tank in the GreenHab. It is nice to be able to water from inside the building and will provide some temperature buffering capacity.

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of health and safety issues: nothing to report.

Notes, questions, concerns and requests to Mission Support: nothing at this time

Copyright © The Mars Society. All rights reserved. | Main Site